Reports until 12:34, Monday 02 December 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:34, Monday 02 December 2024 - last comment - 16:59, Monday 02 December 2024(12024)
B111 Lab Progress

[Torrey, Ian]

General progress notes:

-The three photo diodes required for OFC1 activities are powered.

-All cables to and from the moku/patch panel are set up for OFC1 activities, including PDs, EOM, and DC modulation for the laser.

-1550 path is realigned and can be locked with the laser.

 

Problem:

The ~900 Hz signal that arrises from having a cable plugged into the DC modulation port of the laser needs to be solved. This is a known problem from when we were in B102 but it seems much worse over here. Here is what I know. 

1) Light is blocked. DC modulation port plugged in. ~890 Hz signal observed on PD. This is not optical and must be electrical.

2) Light blocked. DC modulation port is unplugged from the patch panel. Oscillations are gone. To clarify, I am unplugging it here: PXL_20241202_193738276~2.jpg

3) This interferes with the error signal enough to not give a good laser lock. 

4) The oscillations can be seen when the channel is DC coupled and light is hitting the PD.

5) Both the laser and the photo detector are plugged into a blue outlet, but not the same blue outlet.

If anyone has any insights on this let me know. I am curious to see if the PD is plugged into a white outlight if this makes a difference. Lee has suggested a type of circuit to fix this potential grounding issue in the past although I am blanking on the name of it. 

Images attached to this report
Comments related to this report
Torrey Cullen - 16:59, Monday 02 December 2024 (12026)

After some frustrating moku glitches and then some realignment, the second output filter cavity now has 775 light. Pending both 775 paths, and 1550 in OFC2, we should be ready for filter cavity science again.